Show how two phase regions only need one variable to complete define the phase compositions as opposed to single phase fields which need two variables using the Gibb’s phase rule
Answer #1

according to the gibbs shase Rule: - P+f=c+2 Where :- p= no of phase present F= degree of freedom C = component for single co